# Bound States at Semiconductor–Mott Insulator Interfaces

### v2

@@ I Introduction
In navigating the complexities of these distinct material classes, we employ the hierarchy of correlations [33, 34, 33], a framework that aligns both weakly and strongly interacting materials on equal footing while preserving spatial resolution [35]. This approach allows us to bridge the gap between the different description requirements of semiconductors and Mott insulators, offering a unified perspective on their behavior within heterostructures. Within this approach, we will examine the different spin backgrounds of Mott insulators, the unpolarized background and the Mott-Néel state and their influence on charge bound states.
@@ II Hubbard Model and Hierarchy of Correlations
The parameters $U_{\mu}$ and $V_{\mu}$ serve to differentiate between strongly correlated systems ($U\neq 0$, $V=0$) and weakly correlated systems ($U=0$, $V\neq 0$), as the on-site repulsion $U_{\mu}$ and on-site potential $V_{\mu}$ play crucial roles in characterizing the nature of electronic correlations and the relative band alignments within the system. Note that a possibly small $U$ in a weakly correlated semiconductor could be treated by a mean-field approximation and hence simply be absorbed in the magnitude of an effective $V_{\mu}$. This is similar to Fermi liquid theory [37, 38].
@@ II.1 Hierarchy of Correlations
for doublons $I=1$ and holes $I=0$. From these, we define the correlation functions $\langle\hat{c}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\text{corr}}$ [35]. For the relevant part of the dynamics, this can further be simlified by a factorization [43, 39], which yields the doublon $p_{\mu}^{1}$ and holon $p_{\mu}^{0}$ amplitudes as:
@@
The two amplitudes can be grouped together using a spinor notation and governing equations for these (quasi) particles [35] can be derived. In a hyper-cubic lattice dependent on the parallel momentum $k^{\|}$ they read:
@@
The expectation values $\langle\hat{n}_{\mu}^{I}\rangle^{0}$ are to be taken in the mean-field background $\hat{\rho}^{0}_{\mu}$ and we use $U_{\mu}^{I}=IU_{\mu}$. More details on the derivation are presented in App. B.
II.2 Unpolarized Mean-Field Background
@@
This might be created by a finite temperature destroying any spin ordering but to low to excite charges above the Mott gap $U$ or by spin frustration in lattices.
On the other hand, in the weakly correlated semiconductor region, where the quasi-particles are the real electrons, the on-site density matrix takes different forms for the valence and conduction bands: $\hat{\rho}_{\mu}^{0}=\ket{\uparrow\downarrow}_{\mu}\!\bra{\uparrow\downarrow}$ | for the valence band and $\hat{\rho}_{\mu}^{0}=\ket{0}_{\mu}\!\bra{0}$ | for the conduction band. Within each individual region, quasi-particles can be described using the ansatz $p_{\mu}^{I}=\alpha_{I}e^{i\kappa\mu}+\beta_{I}e^{-i\kappa\mu}$ [35], with the proportionality $Ep_{\mu}^{0}=(E-U)p_{\mu}^{1}$ in the Mott insulator. In the semiconducting sites, either one or the other is zero. The corresponding wave numbers are given by:
@@
$r_{\pm}=e^{\pm i\kappa_{\mathrm{Mott}}}\,\quad s_{\pm}=e^{\pm i\kappa_{\mathrm{semi}}}$ (9)
@@
II.3 Bi-Partite Mean-Field Background
At second, there is the antiferromagnetic Mott-Néel state with its checkerboard structure:
@@ II.3 Bi-Partite Mean-Field Background
In this spin configuration, the unit cell has double the size compared to the unpolarized background. This leads to backfolding in the Brillouin zone, giving two solutions: an even and an odd one. Moreover, the bi-partite structure will imprint on the weakly interacting sites.
To get this, we introduce another index $A$ or $B$ for the respective sublattice. In the Mott-Néel state, the coupled equations for doublons and holons thus read:
$Ep_{\mu}^{0_{A}}$ $=-\left[T_{\mathbf{k}}^{\|}p_{\mu}^{1_{B}}+\frac{T}{Z}\left(p_{\mu+1}^{1_{B}}+p_{\mu-1}^{1_{B}}\right)\right]$ (11)
@@
Again, there is a proportionality between the two quasi-particles $p_{\mu}^{1_{B}}=\beta p_{\mu}^{0_{A}}$ with $\beta=\pm\sqrt{\frac{E}{E-U}}$. With the ansatz $p_{\mu}^{1_{B}}=\mathcal{B}\kappa^{\mu}$ and $p_{\mu}^{0_{A}}=\mathcal{A}\kappa^{\mu}$, and we find the eigenmodes:
@@
±$\pm\sqrt{\left[\frac{Z}{2T}\left(\sqrt{E(E-U)}+T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}$
@@
±$\pm\sqrt{\left[\frac{Z}{2T}\left(\sqrt{E(E-U)}-T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}$
with $\kappa_{1}\kappa_{2}=\kappa_{3}\kappa_{4}=1$. The proportionality between them reads
$B_{i}=\frac{1}{U-E}\left[T_{\mathbf{k}}^{\|}+\frac{T}{Z}\left(\kappa_{i}+\kappa_{i}^{-1}\right)\right]A_{i},$ (13)
@@
Without loss of generality, we can extend the bi-partite structure to the semiconducting half-space, now setting $p_{\mu}^{0_{A}}=\pm\alpha p_{\mu}^{0_{B}}$. Together with the ansatz $\lambda^{\mu}$, we find the eigenmodes for $\alpha=+1$ as:
@@
belonging to the ’odd’ solution.
@@ III Unpolarized Mott-Semiconductor interface
Figure 1: Energy schematics of the bound states at the single unpolarized interface. The solid line gives the relevant semiconductor band $V-T_{\mathbf{k}}$ while the dashed line depicts the two Mott bands around $E\approx 0$ and $E\approx U$. The dotted line describes the defining equation for the bound states that are shown as the thick vertical lines outside the bands.
First, our objective is a single unpolarized Mott-semiconductor interface. We consider a hypercubic lattice where the two half-spaces correspond to a Mott insulator with site index $\mu<0$ and a semiconductor with $\mu\geq 0$. In addition to the infinitely extended continuum states [35], bound states, characterized by their spatial decay away from the interface, may occur. For the complex numbers $r_{a}$ and $s_{b}$ in our ansatz, this implies
$p_{\mu}^{0}=\begin{cases}A_{a}\left(r_{a}\right)^{\mu}&\mu<0\text{ with }\left|r_{a}\right|>1\text{ in Mott},\\ B_{b}\left(s_{b}\right)^{\mu}&\mu\geq 0\text{ with }\left|s_{b}\right|<1\text{in semi.},\end{cases}$ (16)
where the indices $a,b$ stand for the $+$ or -- sign.
While we keep the model parameters $T_{\mu\nu},U_{\mu}$ and $V_{\mu}$ constant inside their respective regions, our analysis shows that an additional on-site potential $\Delta V\neq 0$ at the semiconductor interface site $\mu=0$ is a pre-requisite for obtaining interface states. This interface perturbation arises due to local modifications in the epitaxial interfaces of different materials. It might be introduced by a structural relaxation or deformation [45, 46, 47], by the tilting of the perovskite octahedra as in $\mathrm{LaTiO}_{3}/\mathrm{KTaO}_{3}$ [15] or other local modifications [48, 49].
By considering the relation between particles $p_{\mu}^{1}$ and holes $p_{\mu}^{0}$ in the Mott and semiconductor regions, we can establish boundary conditions at the interface using Eq. 6. These relate the amplitudes with the interface perturbation. After some algebra (see App. C for details), the defining equation for bound states at the single unpolarized interface reads
@@
It is important to note that this condition lacks a solution in the absence of an interface perturbation (see App. C.1). The defining equation is exemplarily shown in Fig. 1. The bound state energy solutions are marked. It looks similar to the Newns-Anderson model describing chemisorption or localized magnetic states in metals [50, 51]. Fig. 2 shows one solution to this equation, As expected for a bound state, the probability distribution exhibits localization at the interface – on the lattice site with the interface perturbation – and displays distinct decay constants towards both the Mott insulator and the semiconductor regions.
Figure 2: Bound state $p_{\mu}^{0}$ and quasi particle probability distribution $|p_{\mu}^{0}|^{2}$ at the interface for the unpolarized background with $E=1.16U$. The parameters are $V=1.1U$, $T=0.4U$ and $\Delta V=0.2U$.

### v3

@@ I Introduction
In navigating the complexities of these distinct material classes, we employ the hierarchy of correlations [33, 34, 33], a framework that aligns both weakly and strongly interacting materials on equal footing while preserving spatial resolution [35]. This approach allows us to bridge the gap between the different description requirements of semiconductors and Mott insulators, offering a unified perspective on their behavior within heterostructures. We derive a Schrödinger-like equation for the quasi-particle wave functions beyond the mean-field approximation. Within this approach, we will examine the different spin backgrounds of Mott insulators, the unpolarized background and the Mott-Néel state and their influence on charge bound states.
@@ II Hubbard Model and Hierarchy of Correlations
The parameter $U_{\mu}$ describes the on-site Coulomb interaction in the Fermi-Hubbard model, and it is non-zero only in the Mott insulator. In the following it sets the energy scale. Note that a possibly small $U$ in a weakly correlated semiconductor could be treated by a mean-field approximation and hence simply be absorbed in the magnitude of an effective $V_{\mu}$. This is similar to Fermi liquid theory [37, 38].
In principle, both the Mott insulator and the weakly correlated layer have an on-site potential $V_{\mu}$, as the Hubbard bands do not necessarily need to be centered around $U/2$. But, as only the relative band alignment is crucial, we can set $V_{\mu\,\in\,\mathrm{Mott}}\equiv 0$ and describe the band alignment simply by the offset $V_{\mu\,\in\,\mathrm{semi}}-V_{\mu\,\in\,\mathrm{Mott}}\equiv V$. This incorporates both the material specific parameters of the band alignment and external electric fields into a single value for the band offset. For example, in the $\mathrm{LaTiO}_{3}/\mathrm{SrTiO_{3}}$ interface the specific band alignment depends on the number of grown layers. This would also be dealt with by the offset $V$.
Therefore, the parameters $U_{\mu}$ and $V_{\mu}$ serve to differentiate between strongly correlated systems ($U\neq 0$, $V=0$) and weakly correlated systems ($U=0$, $V\neq 0$), as the on-site repulsion $U_{\mu}$ and on-site potential $V_{\mu}$ play crucial roles in characterizing the nature of electronic correlations and the relative band alignments within the system.
@@ II.1 Hierarchy of Correlations
for doublons $I=1$ and holes $I=0$. These quasi-particles are the physical excitations within the Mott insulator on top of its half-filled background, and thus form a suitable starting point for describing the physics. From these, we define the correlation functions $\langle\hat{c}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\text{corr}}$ [35]. As we are interested in charge bound states, i.e., electrons and doublons/holons, we only take first order correlations into account. To this order, these are the only non-zero correlation functions. This leaves out the higher order correlations describing spin-fluctuations, particle number correlations and doublon-holon correlations, which would in second order act as source terms for the charge mode dynamics. Therefore, we treat the charge modes on top of a fixed mean-field background. This does not include any back-reaction effects of these modes onto the background. Incorporating this would effectively yield a renormalized hopping [43], altering the dispersion relation. The neglected higher order correlations do not alter the dispersion and would solely act as source terms. For $Z\gg 1$ this is a valid approximation. For the relevant part of the dynamics, this can further be simplified by a factorization [44, 39], which yields the doublon $p_{\mu}^{1}$ and holon $p_{\mu}^{0}$ amplitudes as
@@
In a sense, this is the same as writing the entries of the many-body density operator $\hat{\rho}$ in this doublon-holon basis as $\hat{\rho}_{IJ}=\left(p^{I}\right)^{*}p^{J}$ with the wave functions $p^{I}$. The two amplitudes can be grouped together using a spinor notation and governing equations for these (quasi) particles [35] can be derived. The interface in the systems breaks the translational invariance into this one direction, but leaves the other ones intact. Therefore, the parallel momentum $k^{\|}$ is still a conserved quantity. Hence, we decompose the wave functions $p_{\mu}^{I}=\sum_{k^{\|}}p_{\mu}^{I}(k^{\|})e^{ik^{\|}\cdot x_{\mu}^{\|}}$ into their Fourier components. After this, $\mu$ is a scalar index counting the layers parallel to the interface, and for simplicity of notation we do not explicitly write down the momentum dependence. In a hyper-cubic lattice dependent on the parallel momentum $k^{\|}$ the Schrödinger-like equations for the doublons $I=1$ and holons $I=0$ can be combined using $U_{\mu}^{I}=IU_{\mu}$,
@@
In this eq., $T_{\mathbf{k}}^{\|}=2T/Z\sum_{i}\cos(k_{i}^{\parallel})$ gives the kinetic energy contribution parallel to the interface. Because of the periodicity in this direction bands form. The expectation values $\langle\hat{n}_{\mu}^{I}\rangle^{0}$ are to be taken in the mean-field background $\hat{\rho}^{0}_{\mu}$ and encode the spin background structure. More details on the derivation are presented in App. B. Essentially, the hierarchy of correlations yields two coupled equations for the wave functions of doublons $p_{\mu}^{1}$ and holons $p_{\mu}^{0}$ on top of a half filled background discretized on the lattice. These can be solved for plane wave eigenmodes, evanescent solutions and boundary conditions, as known from the usual Schrödinger equation in quantum mechanics. We might formally write the system as
$\mathcal{H}_{\mu-1}\begin{bmatrix}p_{\mu-1}^{0}\\ p_{\mu-1}^{1}\end{bmatrix}+\mathcal{H}_{\mu+1}\begin{bmatrix}p_{\mu+1}^{0}\\ p_{\mu+1}^{1}\end{bmatrix}=\mathcal{H}_{\mu}\begin{bmatrix}p_{\mu}^{0}\\ p_{\mu}^{1}\end{bmatrix}$. (7)
This is the well-known form of a second-order differential equation in space, such as the Schrödinger , Dirac or Klein-Gordon equation, for wave functions discretized on a lattice.
II.2 Comparison to other Methods
The hierarchy of correlations is based on a formal expansion into the inverse coordination number $1/Z$, not on a small parameter $T/U$ or $U/Z$ as in perturbation theory. This allows us to treat weakly and strongly correlated materials on the same level of theory. The dispersion relation within the strongly correlated Mott insulator might be calculated by means of other methods [45, 46, 47, 48, 49] like the Hubbard-I approximation, Roth’s two-pole scheme, the Second-Order Decoupling Approximation or Composite Operator Methods. Similar to the hierarchy of correlations, these methods rely on truncating an infinite series of expectation values or correlation functions. However, the hierarchy does have a small parameter $1/Z$, leading to a clear separation of leading and sub-leading contributions [50, 51] controlling this truncation, while the other methods do not (see, e.g., [48], Eq. 56), making them inherently uncontrolled. Moreover, while the former approaches are best suited to thermal states, our framework naturally extends to dynamics (see [52] for the thermal case within the hierarchy). The hierarchy leads to non-perturbative results in $T/U$, as there are non-polynomial dependencies [43].
An alternative method that has been used to calculate transport across Mott insulating layers [53] and interface charge order [54, 55] is Dynamical Mean-Field Theory (DMFT) [56]. The guiding idea behind DMFT is different from our approach: DMFT aims at approximating the self-energy of the system in order to find the system’s Green’s function, while the hierarchy aims at finding two- or multi-site correlation functions. The lifetime broadening related to self-energy is only included in higher orders. While in our approach the hopping scales as $1/Z$, DMFT uses a $1/\sqrt{Z}$ scaling. As a consequence, the $Z\to\infty$ limit yields already a non-trivial and physically interesting result in DMFT, while in the hierarchy of correlations all correlations vanish in this limit. This simplicity allows us to calculate the first-order $1/Z$ correlations on top of the mean-field background, and this includes the effects of the lattice structure and dimensionality. The hierarchy of correlations fully accounts for space-time dependence, particularly in higher dimensions [52, 57], while DMFT [56] maps to an effective single lattice site. This single-site mapping leads to a purely local self-energy, neglecting all non-local correlations. DMFT is formally exact only in the limit of infinite dimensions, and corrections beyond this limit are generally uncontrolled. It is known to fail in low dimensions at low temperatures. In contrast, our approach incorporates non-local correlations, and the validity of the hierarchy of correlations is not dependent on the dimensionality of the system, but only on the number of neighbors. Moreover, the resulting equations are simple enough for analytic treatment, whereas DMFT is primarily numerical [58, 59].
The hierarchy of correlations is expected to fail for low coordination numbers, as terms $1/Z^{2}$ cannot be neglected compared to $1/Z$ anymore, but even in these cases the results are still giving qualitative and quantitative agreement after including second order back-reaction effects and renormalizing the hopping strength, see Ref. [57] for a comparison with exact diagonalization. Another situation where the method is expected to break down occurs when the two-point correlations become comparable in magnitude to the on-site expectation values. This typically happens near a phase transition, for instance during a quench from the Mott insulating to the superfluid phase in the Bose–Hubbard model. In such a regime, one would need to describe the evolution of the correlations on top of a newly determined mean-field background. Moreover, the hierarchy either works in the weakly or strongly interacting regime, but cannot work in the intermediate interacting strength relevant for, e.g., Kondo physics. DMFT, on the other hand, is capable of doing both things [54].
To demonstrate the applicability, we assume a hyper-cubic lattice mostly for simplicity, though some relevant materials are cubic indeed. Many perovskite Mott insulators, such as TiF3 [60] and LaTiO3 [29] exhibit insulating behavior on three-dimensional cubic lattices with $Z=6$. Others occur on (quasi-)two-dimensional triangular lattices with the same coordination number [61, 62, 63]. Since the hierarchy depends on coordination number rather than dimensionality, our scaling applies equally to both cases. Second-order effects can also be included via a renormalized hopping parameter [43], ensuring robustness across these systems. At present, such higher-order calculations exist only for homogeneous systems [64] and remain a subject for future work.
II.3 Parameter Choice
Even though we describe the complex interactions in the Mott insulator using a simplified Fermi–Hubbard model, there are physically well-motivated choices for the parameters. A common strategy is to use DFT calculations as a complementary tool to obtain the band structure of the material of interest, then fit a tight-binding Hamiltonian to extract effective parameters. This approach generally yields a hopping-to-interaction ratio in the range $0.05<(T/Z)/U<0.2$ [54, 53, 65, 66, 67].
For instance, the DMFT study in Ref. [68] examined a heterostructure with a finite number of Mott-insulating layers embedded in an infinite band insulator. They used parameters $T=0.3\,\mathrm{eV}$ and $U=4.8\,\mathrm{eV}$, values extracted from experimental work on $\mathrm{SrTiO}_{3}/\mathrm{LaTiO}_{3}$ superlattices [1], corresponding to $T/U=0.0625$. In Ref. [53], DMFT combined with the Keldysh formalism was employed to study a Mott insulator coupled to metallic leads, using $T/U=0.066$ and a band offset of $V=U/2$. Similarly, Ref. [55] investigated a Mott–metal interface with a band offset of approximately $V\approx 0.9U$.
Experimentally, band offsets have been reported in a range of systems: about $V=0.05U$ for Ni–NiO–Ni junctions, $V=0.3U$ for Ni–MnO–Ni junctions [69], $V\approx 0.55U$ for a $\mathrm{LaVO}_{3}/\mathrm{SrTiO}_{3}$ interface [70], and $V\approx 0.3U$ and $V\approx 0.25U$ for the conduction band offsets in $\mathrm{SrTiO}_{3}/\mathrm{SmTiO}_{3}$ and $\mathrm{SrTiO}_{3}/\mathrm{GdTiO}_{3}$, respectively [71].
In the following, we will use either $T=0.2U$ or $T=0.4U$ as the hopping strength. This yields an effective strength of $T/Z=0.05U$ to $T/Z=0.1U$, right in the range of other studies. We checked that varying the parameters within this range leads to only quantitative, but not qualitative changes, as the parameters remain within the (gapped) Mott-Hubbard regime.
II.4 Unpolarized Mean-Field Background
@@
[…]
On the other hand, in the weakly correlated semiconductor region, where the quasi-particles are the real electrons, the on-site density matrix takes different forms for the valence and conduction bands: $|\hat{\rho}_{\mu}^{0}=\ket{\uparrow\downarrow}_{\mu}\!\bra{\uparrow\downarrow}$ for the valence band and $|\hat{\rho}_{\mu}^{0}=\ket{0}_{\mu}\!\bra{0}$ for the conduction band.
Within each individual region, quasi-particles eigenmodes can be described using the ansatz $p_{\mu}^{I}=\alpha_{I}e^{i\kappa\mu}+\beta_{I}e^{-i\kappa\mu}$ [35], with the proportionality $Ep_{\mu}^{0}=(E-U)p_{\mu}^{1}$ in the Mott insulator. In the semiconducting sites, either one or the other is zero. These describe the plane wave eigenmodes of the coupled doublons and holons in the Mott insulator and the electrons in the semiconductor, respectively. The corresponding wave numbers are given by
@@
$r_{\pm}=e^{\pm i\kappa_{\mathrm{Mott}}}\,,\quad s_{\pm}=e^{\pm i\kappa_{\mathrm{semi}}}$ (10)
@@
The quasi-particles (holons and doublons) show a wave like behavior inside the Mott bands and might be described by plane waves, similar to the electrons in the semiconducting band. The wave functions of these quasi-particles do not evolve independent of each other, but they are coupled.
II.5 Bi-Partite Mean-Field Background
Secondly, there is the antiferromagnetic Mott-Néel state with its checkerboard structure:
@@ II.5 Bi-Partite Mean-Field Background
There are two different sublattices $A$ and $B$ with either spin-up or spin-down electrons. Without loss of generality, we use $\langle\hat{n}_{\mu_{A}\uparrow}\rangle=1$, $\langle\hat{n}_{\mu_{B}\uparrow}\rangle=0$. In this spin configuration, the unit cell has double the size compared to the unpolarized background. This leads to backfolding in the Brillouin zone, giving two solutions: an even and an odd one. Moreover, the bi-partite structure will imprint on the weakly interacting sites.
In order to capture the quasi-particles and their eigenmodes on this bi-partite lattice structure, we introduce another index $A$ or $B$ for the respective sublattice. In the Mott-Néel state, the coupled equations for doublons and holons thus read
$Ep_{\mu}^{0_{A}}$ $=-\left[T_{\mathbf{k}}^{\|}p_{\mu}^{1_{B}}+\frac{T}{Z}\left(p_{\mu+1}^{1_{B}}+p_{\mu-1}^{1_{B}}\right)\right],$ (12)
@@
As in the unpolarized state, doublons and holons are not independent of each other, but they are coupled. This leads to a proportionality between the two quasi-particles $p_{\mu}^{1_{B}}=\beta p_{\mu}^{0_{A}}$ with $\beta=\pm\sqrt{\frac{E}{E-U}}$. With the ansatz $p_{\mu}^{1_{B}}=\mathcal{B}\kappa^{\mu}$ and $p_{\mu}^{0_{A}}=\mathcal{A}\kappa^{\mu}$, $\mathcal{A}$ and $\mathcal{B}$ being the wave function amplitudes, we find the eigenmodes
@@
$\pm\sqrt{\left[\frac{Z}{2T}\left(\sqrt{E(E-U)}+T_{\mathbf{k}}^{\|}\right)\right]^{2}-1},$
@@
$\pm\sqrt{\left[\frac{Z}{2T}\left(\sqrt{E(E-U)}-T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}$
with $\kappa_{1}\kappa_{2}=\kappa_{3}\kappa_{4}=1$. Because of the backfolding there are four eigenmodes. Within the Hubbard bands, these obey $|\kappa_{i}|=1$, such that we have plane waves again. This can be seen from the identity $x\pm i\sqrt{1-x^{2}}=e^{\pm i\arccos(x)}$, which defines the wave numbers for the two different types of plane wave eigenmodes $\kappa_{1,2}$ and $\kappa_{3,4}$. Within the Mott bands, they read:
$\cos(x_{1,2})$ $=\frac{Z}{2T}\left(\sqrt{E(E-U)}+T_{\mathbf{k}}^{\|}\right),$ (14)
$\cos(x_{3,4})$ $=\frac{Z}{2T}\left(\sqrt{E(E-U)}-T_{\mathbf{k}}^{\|}\right).$
Outside the Mott bands, they define decaying solutions. The proportionality between them reads
$\mathcal{B}_{i}=\frac{1}{U-E}\left[T_{\mathbf{k}}^{\|}+\frac{T}{Z}\left(\kappa_{i}+\kappa_{i}^{-1}\right)\right]\mathcal{A}_{i},$ (15)
@@
Without loss of generality, we can extend the bi-partite structure to the semiconducting half-space, now setting $p_{\mu}^{0_{A}}=\pm\alpha p_{\mu}^{0_{B}}$ in the conduction band case. The eigenmodes on the two sublattices are not independent of each other. Together with the ansatz $\lambda^{\mu}$, we find the eigenmodes for $\alpha=+1$ as:
@@
belonging to the ’odd’ solution. Inside the semiconducting band these are plane waves $e^{ik}$ with wave numbers defined as solution of $\cos(k)=\frac{Z}{2T}(E-V+T^{\|}_{\mathbf{k}})$ and $\cos(k)=\frac{Z}{2T}(E-V-T^{\|}_{\mathbf{k}})$, respectively.
As in the unpolarized case, the quasi-particles on the bi-partite lattice show a wave-like behavior within the bands. They are coupled, and the wave functions on the different sublattices are proportional to each other with an energy dependent factor. Moreover, there are two different solutions, even and odd ones, with the same sign or a sign switch between the wave functions on their respective sublattice.
@@ III Unpolarized Mott-Semiconductor interface
Figure 1: Density of states $D(E)$ of the Mott insulator (blue dashed) and the semiconductor (red solid) together with the delta-peaks of the bound states shown as the black dashed vertical lines.
First, our objective is a single unpolarized Mott-semiconductor interface. We consider a hypercubic lattice where the two half-spaces correspond to a Mott insulator with site index $\mu<0$ and a semiconductor with $\mu\geq 0$. In addition to the infinitely extended continuum states [35], bound states, characterized by their spatial decay away from the interface, may occur. For the quasi-particle wave functions $r_{a}$ and $s_{b}$ in our ansatz, this implies
[…]
where the indices $a,b$ stand for the ++ or - sign in Eq. 10. Outside their respective bands, both $\kappa_{\mathrm{Mott}}$ and $\kappa_{\mathrm{semi}}$ are purely imaginary, such that the holon wave function $p_{\mu}^{0}$ decays away from the interface to both sites. The bound state energy finally decides if the physically sound solution is comprised of the ++ or - sign solution.
While we keep the model parameters $T_{\mu\nu},U_{\mu}$ and $V_{\mu}$ constant inside their respective regions, our analysis shows that an additional on-site perturbation $\Delta V\neq 0$ at the semiconductor interface site $\mu=0$ is a pre-requisite for obtaining interface states. This interface perturbation arises due to local modifications in the epitaxial interfaces of different materials. It might be introduced by a structural relaxation or deformation [73, 66, 74], by the tilting of the perovskite octahedra as in $\mathrm{LaTiO}_{3}/\mathrm{KTaO}_{3}$ [15] or other local modifications [75, 76].
By considering the relation between doublons $p_{\mu}^{1}$ and holons $p_{\mu}^{0}$ in the Mott and semiconductor regions, we can establish boundary conditions at the interface using Eq. 6. These relate the amplitudes with the interface perturbation. After some algebra (see App. C for details), the defining equation for bound states at the single unpolarized interface reads:
@@
It is important to note that this condition lacks a solution in the absence of an interface perturbation (see App. C.1). Without it, the interface only supports interface resonances, decaying to one site while being a plane wave in the other half-space. Fig. 1 shows the density of states of this system as the bands of the Mott and semiconductor together with the delta-like bound state energies. It looks similar to the Newns-Anderson model describing chemisorption or localized magnetic states in metals [77, 78]. Fig. 2 shows one solution to this equation. As expected for a bound state, the probability distribution exhibits localization at the interface – on the lattice site with the interface perturbation – and displays distinct decay constants towards both the Mott insulator and the semiconductor regions.
We find bound states solutions for attractive and repulsive interface perturbations. This is a distinction from the usual quantum mechanics case with a delta potential, in which only the attractive case has bound state solutions.
Figure 2: Bound state for the holon wave function $p_{\mu}^{0}$ and quasi particle probability distribution $|p_{\mu}^{0}|^{2}$ at the interface for the unpolarized background with $E=1.16U$. The parameters are $V=1.1U$, $T=0.4U$ and $\Delta V=0.2U$.